There was a problem with filling 1L12. The filling rate was very low. Troubleshooting required tunnel access. During this access adjustments were made to the JT LVDT. Also the RT was not fully open. After these adjustments, the liquid level fill rate has drastically improved.
In addition to that, the primary supply u-tube is sweaty and will need to be pumped at the next maintenance day.
The fast charge module in the NL03 vacuum crate was found to be defective.
After securing the MOPS for the vacuum crate and disconnecting all slow valve
cables I replaced the defective card. I then reconnected cables for the two
adjacent valves needed to ready the fast valve and re-energized the crate.
The fast valve opened successfully with a good "ready" indicator. Once the
reaming valves were reconnected and opened, the "Beam Permit" light on the
FSD card came came on.
FSD is now clear!
Gary Croke checked the recent attempts over the past several days to set the 0L06 spectrometer dipole to straight ahead mode. Going to straight ahead mode requires the magnet to switch polarity. He verified that the polarity switch command is being issued but the magnet power supply is not confirming the switch. The ioc waited as long as ~45 seconds for the confirmation before giving up.